A survey of

2287

adults in a certain large country aged 18 and older conducted by a reputable polling organization found that

406

have donated blood in the past two years. Complete parts​ (a) through​ (c) below.

Click here to view the standard normal distribution table (page 1).

LOADING...

Click here to view the standard normal distribution table (page 2).

LOADING...

​(a) Obtain a point estimate for the population proportion of adults in the country aged 18 and older who have donated blood in the past two years.

ModifyingAbove p with caret

equals. 178

​(Round to three decimal places as​ needed.)

​(b) Verify that the requirements for constructing a confidence interval about p are satisfied.

The sample

can be assumed to be

cannot be assumed to be

is stated to be

is stated to not be

a simple random​ sample, the value of

n ModifyingAbove p with caret

ModifyingAbove p with caret left parenthesis 1 minus ModifyingAbove p with caret right parenthesis n ModifyingAbove p with caret left parenthesis 1 minus ModifyingAbove p with caret right parenthesis

n

ModifyingAbove p with caret

is

nothing

​,

which is

greater than or equal to

less than

​10, and the

sample size

sample proportion

population size

population proportion

is stated to not be

is stated to be

cannot be assumed to be

can be assumed to be

less than or equal to​ 5% of the

sample size.

population size.

sample proportion.

population proportion.

​(Round to three decimal places as​ needed.)
Answer #1

a) The point estimate is = 406/2287 = 0.178

B) The sample a simple random sample, the value of is greater than Or equal to 10 and the sample size is less than or equal to 5% of the population size.

= 2287 * 0.178 * (1 - 0.178) = 334.625